OpenStudy (anonymous):

tan(pix^2/4) differentiate

OpenStudy (anonymous):

\[\tan(pix^2/4)\]

OpenStudy (sheraz12345):

\[\sec ^{2}\left(\begin{matrix}pix ^{2} \\ 4\end{matrix}\right)(2x \frac{ \pi }{ 4 })\]